may15:Alfa calibration scans, rotation angle


Compare to may14 results after re-installation

    Alfa was reinstalled in the dome on 28apr15  after being worked on in the lab. Spider scans (heiles calibration scans) were taken with alfa on 30apr15, 01may15 using beams0..6
This data was taken with model18A  and with the tiedowns active.
The 30apr15 data had not signal in beam 6b (the switch for the filter was broken) so beam 6 data from this set was discarded. The problem was fixed for the 01may15 data.

    The hf screen had been installed:

The setup:

Calibration run results:

The plots show the   calibration results (.ps) (.pdf):

Fitting for rotation angle error:

    Part of the pointing error could come from the alfa rotation angle being off. To test for this :
    Plots showing the pointing error vs rotation angle offset (.ps) (.pdf):


processing: 150501/